Sum of a certain number and a second number is 24. The first number minus the second is 40.
Answer by mananth(16946)

let the numbers be x & y
x is greater than y
x+y=24
x-y=40
add the two equations
2x=64
/2
x=32
plug x in the equation x+y=24
32+y=24
y=24-32
y=-8